An L-C circuit has an inductance of 0.440 H and a capacitance of 0.220 nF . During the current oscillations, the maximum current in the inductor is 1.00 A .

Part A:

What is the maximum energy Emax stored in the capacitor at any time during the current oscillations? Express your answer in joules.

Part B:

How many times per second does the capacitor contain the amount of energy found in part A? Express your answer in times per second.
Answer #1

using conservation of energy

max capacitive energy = max inductive energy

0.5 CV^2 = 0.5 Li^2

E = 0.5* 0.44* 1^2 = 0.22 J

=======

number of timesit direction changes

N = 2 * oscillating frequency

N = 2* 1 /( 2 pi  \sqrt( CL))

N = 32369

=======

Comment in case any doubt, will reply for sure.. Goodluck